Probability and Statistics I GSIA, Carnegie Mellon University

45-733, Spring 2002

Homework #2, due Thursday, January 31, 2002

Please do the following problems from the book:

pp 141-4 3,5,11,13
pp 154-6 19,23
pp 171-7 29,33,49,59

If you get stuck, it would be wise to try to work even-numbered problems similar to the ones assigned, since many of these have solutions in the back of the book.